Solve each system of equations. Check the solutions.
2x-y=4
x+2y=7

Answer by checkley75(3666)

2x-y=4
x+2y=7 or x=7-2y
2(7-2y)-y=4
14-4y-y=4
-5y=4-14
-5y=-10
y=-10/-5
y=2 ans.
2x+2=4
2x=4+2
2x=6
x=6/2
x=3 ans.
Proof:
3+2*2=7
3+4=7
7=7